OpenStudy (anonymous):

The length of a rectangle is 5 times the width. What is the perimeter of the rectangle if the width is 12 inches? A. 72 inches B. 120 inches C. 144 inches D. 240 inches

OpenStudy (anonymous):

help

OpenStudy (confluxepic):

l=5w w=12 l=5(12) l=60

OpenStudy (confluxepic):

P=2(l+w) P=2(60+12) P=2(72) P=144

OpenStudy (confluxepic):

The answer is C.

OpenStudy (anonymous):

thx

OpenStudy (confluxepic):

@schoolfamily101

OpenStudy (anonymous):

ok thx

OpenStudy (confluxepic):

I know I'm one year late but I keep getting notifications from this post. \(\color{blue}{\text{Originally Posted by}}\) @schoolfamily101 @~ ok thx \(\color{blue}{\text{End of Quote}}\) You're welcome. :)